OpenStudy (anonymous):

Solve simultaneously for "x" and "y" 2x + 5y = 28 x - 2y = 2

OpenStudy (anonymous):

2x + 5y = 28 2x - 4y = 4 subtract 9y = 24 y = 24/9 = 8/3 sub y = 8/3 in x - 2y = 2 x - 2*8/3 = 2 x = 16/3 + 2 = 22/3 roots are x = 22/3 , y = 8/3